OpenStudy (anonymous):

divide and simplify. y^2-1/y divided by y+1/y+4

OpenStudy (anonymous):

Rewrite \(y^2-1\) using the difference of squares formula: \[a^2 - b^2 = (a+b)(a-b)\]

OpenStudy (anonymous):

what do u mean?

OpenStudy (anonymous):

\[y^2 - 1\] is the difference of \(y^2\) and \(1^2=1\)

OpenStudy (anonymous):

no, its exactly as i typed

OpenStudy (anonymous):

You missunderstand. \[y^2 - 1 = y^2 - 1^2\] since \(1 = 1^2\)

OpenStudy (anonymous):

And if you have anything like \[a^2 - b^2\] You can write it as \[(a+b)(a-b)\]